determine the amount cu2s at 85% we need to obtain 1 kg of cuobif the yield of tje reaction is 92%
1
Expert's answer
2019-12-22T02:38:18-0500

Cu2S + 2 O2 → 2 CuO + SO2

m(Cu2S)=M(Cu2S)*m(CuO)/M(CuO)=159.16(g/mol)*1000(g)/79.54(g/mol)=2001(g)

2001g/0.92=2175g

2 175g/0.85=2558.8g

Answer:2558.8g